The company XYZ’s next year dividend per share is expected to be 4.50. The dividend in subsequent years is expected to grow at a rate of 10% per year. If the required rate of return is 15% per year, what should be its price? The prevailing market price is 80.

1
Expert's answer
2021-10-31T18:26:52-0400

price:


V0 = D1 / (Ke - g ) ;


D1 = Expected dividend next year

g = Expected growth rate in dividends

Ke = Investor’s required rate of return


"V_0=\frac{4.5}{0.15-0.10}=90"